OpenStudy (anonymous):

LAST ONE :D Solve x – 2y = 5 using the substitution method. 2x – 4y = 1

OpenStudy (lgbasallote):

10 -8y = 1 y = -9/8 x =

OpenStudy (lgbasallote):

haha...i cant think with all this pressure @_@

OpenStudy (bahrom7893):

no pressure, im not answering this time..

OpenStudy (anonymous):

I wonder why so much attention? Is it because she is giving love and medals both? :P

OpenStudy (lgbasallote):

hershey is getting ready...

OpenStudy (anonymous):

answer to yo question :P 10 - 8y = 1 , y = -9/8 , x= which igd won -_- ...

OpenStudy (anonymous):

all i know right meow. is im eating a good as pellet grilled cheese.

OpenStudy (anonymous):

@ FoolsforMath , shes pretty :D lol.. and nicest one in here.

OpenStudy (anonymous):

haha well said bahroom :D

OpenStudy (anonymous):

lol bahrom, I don't' think she was serious when she asked you to marry her :P

OpenStudy (anonymous):

Thanks Hershey :3 You're the nicest one ive met so far also. And i can post 623687597645 more pictures if youd like. facebook me, crazies.

OpenStudy (bahrom7893):

lol obviously

OpenStudy (anonymous):

Your welcome. And thanks -^^ Lol

OpenStudy (anonymous):

623687597645 -- can you write this in words?

OpenStudy (lgbasallote):

aww..whyd you have to break the news Fools...we all would've liked to live in our fantasies that she was serious a little while longer :(

OpenStudy (anonymous):

But really all you thanks for helping me. I really do appreciate it.

OpenStudy (anonymous):

aww I am sorry igb :D

OpenStudy (anonymous):

oh lol :D lgb

OpenStudy (anonymous):

@ilikegoosmells no need to thank us. We help anyways :D

OpenStudy (anonymous):

Most educated people ive ever talked to over the internet :p

OpenStudy (anonymous):

<3

OpenStudy (lgbasallote):

...and funniest group :))) this is the funniest conversation i had in this site =))))

OpenStudy (anonymous):

It's all get funny with cute girls around.

OpenStudy (lgbasallote):

@Fool too bad i gave you a medal already...that was an epic comment :))

OpenStudy (anonymous):

help me with my question D; lol..

OpenStudy (anonymous):

Haha gooooooood. im a lil upset right meow though...

OpenStudy (anonymous):

my grilled cheese is gone.

OpenStudy (anonymous):

haha Thanks lgb, but I have to go now. Dinner.. see you all around :)

OpenStudy (lgbasallote):

im not new here bro =))) ive been her like 5 months..i just dont have laughs like this =))) but good for you you got the chance to use it

OpenStudy (anonymous):

I wish i could give you guys 83979482 medals

OpenStudy (lgbasallote):

you are...:p hahaha im probably 10 years younger than you haha justkidding

OpenStudy (lgbasallote):

random numbers are random :))

OpenStudy (anonymous):

Im probably youngest. Only 16 :l

OpenStudy (lgbasallote):

don't worry ilike...im just 16 too :) turning 17 this May. You are not alone :D

OpenStudy (lgbasallote):

why? how old are you?

OpenStudy (lgbasallote):

frankly...yes i thought so :)) there are many people here who are old already =)) how old are you?